Comida corp. is a grocery store located in the midwest. It paid an annual dividend of ​$3.00 last year to its shareholders and plans to increase the dividend annually at 4.0​% forever. It has 400,000 shares outstanding. The shares currently sell for $21.25 per share. Comida Corp. has 10,000 semiannual bonds outstanding with a coupon rate of 8​%, a maturity of 18 years, and a par value of​$1,000. The bonds currently have a yield to maturity​ (YTM) of 7​% per bond. What is the adjusted WACC for Comida Corp. if the corporate tax rate is 35​%?

What is the adjusted WACC for Comida corp. if the corporate tax rate is 35​%? ___​%

​(Round to two decimal​ places.)
Answer #1

Cost of Equity: Current Dividend = D0 = $ 3, Perpetual Growth Rate = g = 4%, Current Price = P0 = $ 21.25

Expected Dividend = D1 = D0 x (1+g) = 3 x (1.04) = $ 3.12

Cost of Equity = ke = (D1/P0) + g = (3.12/21.25) + 0.04 = 0.18682 or 18.682 % ~ 18.68 %

Number of Common Stock = 400000,

Market Value of Equity = 400000 x 21.25 = $ 8500000 = E

Cost of Debt: Number of Bonds Issued = 10000, Coupon Rate = 8%, Compounding Frequency:S emi-Annual, Maturity = 18 years or (18 x 2) = 36 semi-annual periods, Yield to Maturity (YTM) = kd = 7%, Par Value = $ 1000

Semi-Annual Coupon = 0.08 x 1000 x 0.5 = $ 40

Bond Price = 40 x (1/0.035) x [1-{1/(1.035)^(36)}] + 1000 / (1.035)^(36) = $ 1101.45

Market Value of Debt = D = 1101.45 x 10000 = $ 11014524.7 = D

Total Market Value = T = D + E = 8500000 + 11014524.7 = $ 19514524.7

Equity Proportion = e = (8500000/19514524.7) = 0.4356 and Debt Proportion = d = 11014524.7 / 19514524.7 = 0.5644

Tax Rate = t = 35 %

WACC = e x ke + (1-t) x kd x d = 0.4356 x 18.68+ 0.5644 x (1-0.35) x 7 = 10.7046 % ~ 10.70